OpenStudy (anonymous):

7 square root 2 multiply 4 square root 5 please help and explain

OpenStudy (anonymous):

28square root 10

OpenStudy (anonymous):

\[7\sqrt{2}*4\sqrt{5}\] 7*4=28 \[\sqrt2*\sqrt{5}=\sqrt{10}\] so, \[7\sqrt{2}*4\sqrt{5}=28\sqrt{10}\]